simplify( (x-y)/(z)+(y-z)/(x)+(z-x)/(y) )( (z)/(x-y)+(x)/(y-z)+(y)/(z-x) )if x+y+z=o

Grade:9

2 Answers

Pavan Kumar
6 years ago
Dear student,
thank you for asking the question here is the solution
let x+y = a . so we can wrote x=a cos2¢,y=a  sin2¢, then z= -a,
substitute the above values in the above equation
(x-y)/z= (a cos2¢-a  sin2¢)/a= cos2¢- sin2¢
(y-z)/(x)=(a  sin2¢-a)/(a cos2¢)=   -1 
(z-x)/(y) =(a-a cos2¢)/a  sin2¢=1
then the above equation becomes
(cos2¢- sin2¢)(1/cos2¢- sin2¢)=1
the answer is 1,
if you have any doubts remaining let me know
Pavan Kumar
6 years ago
Dear student there is something to remind you
those are,
cos2,sin2.
dont thing in another way.
sorry for that typing mistake.